f 0f g {3} when f=2x and g = 5x-1
Answer by stanbon(75887)

f 0f g {3} when f=2x and g = 5x-1
--------------
fog(3) = f[g(3)] = f[5*3-1] = f[14] = 2*14 = 28
=====================================================
Cheers,
Stan H.